a nurse wants to add water to 30 ounces of a 10% solution of benzakonium chloride to dilute it to an 8% solution how much water must she add?

I'm not sure if I doing it right
8(x+30) + 0.10x

Found 2 solutions by mananth, lwsshak3:
Answer by mananth(16946)

percent ---------------- quantity
solution10 ---------------- 30
water 0 ---------------- x
Total 8 ---------------- 30+x

10*30+0*x=8(30+x)

300+0 =240+8x

-8x=240 -300

-8x=-60

/-8

x=7.5 liters of water

Answer by lwsshak3(11628)

a nurse wants to add water to 30 ounces of a 10% solution of benzakonium chloride to dilute it to an 8% solution how much water must she add?
..
letx=ounces of water to add
then 30+x=ounces of diluted solution
remember, when you add water, you are adding 0% solution
10%(30)+0%(x)=8%(30+x)
3=2.4+.08x
.08x=3-2.4=.6
x=7.5 ounces
ans:
The nurse must add 7.5 ounces of water to dilute to an 8% solution